a bag contains 5 red marbles and 10 blue marbles. what is the probability that the two marbles drawn from the bag (randomly without replacement) will both be blue?

Answer by edjones(8007)

5+10=15
10/15 * 9/14 = 90/210 = 3/7
.
Ed